# Exact WKB in all sectors I: Potentials with degenerate saddles

### v1

@@ 2.3 Dictionary
Let us finally discuss the connection between Airy and Weber type EWKB in a more quantitative manner and introduce a dictionary between the expressions of the WKB actions in both approaches. Note that this was originally discussed in Sueishi:2021xti ; DDP2 , but our construction covers more general cases and stating the explicit dictionary serves as a guideline in later sections.
To make a direct connection to the literature which uses the all-order WKB expansion (e.g. see Basar:2017hpr ; Basar:2015xna ), let us recall the canonical definitions of the WKB action and its dual as
@@
Then, recalling the construction in Sueishi:2021xti , we put forward a dictionary between Airy- and Weber-type cycles as follows:
@@
      $\Pi_{A}$ $=e^{-\frac{2\pi i}{g}a(u,g)}=e^{-\mathcal{G}^{\mathrm{D}}(i\frac{u}{\hbar},ig)}\,\frac{2\pi\left(\frac{g}{\mathcal{C}_{\mathrm{i}}}\right)^{-2\mathcal{F}^{\mathrm{D}}(i\frac{u}{g},ig)}}{\left[\Gamma\left(\frac{1}{2}+\mathcal{F}^{\mathrm{D}}\left(i\frac{u}{g},ig\right)\right)\right]^{2}}\,,\quad\,(u_{0}=u_{\mathrm{T}})\,.$ (88)
@@
      $\Pi_{A}$ $=e^{-\frac{2\pi i}{g}a(u,g)}=e^{-\mathcal{G}^{\mathrm{D}}(i\frac{u}{\hbar},ig)}\,\frac{\sqrt{2\pi}\left(\frac{g}{\mathcal{C}_{\mathrm{ii}}}\right)^{-\mathcal{F}^{\mathrm{D}}(i\frac{u}{g},ig)}}{\Gamma\left(\frac{1}{2}+\mathcal{F}^{\mathrm{D}}\left(i\frac{u}{g},ig\right)\right)}\,,\qquad(u_{0}=u_{\mathrm{T}})\,.$ (90)
@@
Note that we introduce $\mathcal{F}^{\mathrm{D}}$ and $\mathcal{G}^{\mathrm{D}}$ in the expressions (86), (88) and (90). They refer to the $S$-dual case characterized by $V_{\mathrm{D}}=u_{\mathrm{T}}-V$. The appearing imaginary factors are related to the inversion of a well. This is the situation when we compute the functions $\mathcal{F}$ and $\mathcal{G}$ around a barrier top which are related to the dual functions as
@@ 3 Chebyshev potentials, transition and duality
    We are elaborating in their discussion with a focus on the transition from $u<1$ sector to $u>1$ sector, our findings in this section provide guidelines for the later sections where we analyze periodic and symmetric double-well potentials.
@@
    We also provide a conjecture on an $S$-duality transformation of the well-known P-NP relation for genus-1 potentials in sector $u<1$.
@@
A resurgence relation between perturbative and non-perturbative physics can be expressed by the WKB actions $a(u,\hbar)$ and $a^{\mathrm{D}}(u,\hbar)$ or equivalently $\mathcal{F}(\tilde{u},\hbar)$ and $\mathcal{G}(\tilde{u},\hbar)$. In the EWKB framework, it is more natural to obtain a relation between $\mathcal{F}$ and $\mathcal{G}$: Around the bottom of wells and for genus-1 potentials [^7] More specifically, (92) is for the cases when all the actions of the wells are proportional to each others which indicates there are only two independent WKB actions. Recently, an generalization was presented in Cavusoglu:2023bai ; Cavusoglu:2024usn . The authors showed a paramteric form P-NP holds for the deformed quartic potentials which reduces to cubic and quartic cases in extreme limits and P-NP relation takes the form of (92)., this relationship is well-known in various forms Alvarez1 ; Alvarez2 ; Alvarez3 ; Dunne:2013ada ; Dunne:2014bca and in general, it can be expressed as
@@
$\omega_{0}^{\mathrm{D}}(u)\,a_{0}(u)-a_{0}^{\mathrm{D}}(u)\,\omega_{0}(u)=2ia_{0}^{\mathrm{D}}(0)\,\omega_{0}(0)\,,$ (93)
where $\omega_{0}=\frac{{\mathrm{d}}a_{0}}{{\mathrm{d}}u}$ and $\omega_{0}^{\mathrm{D}}=\frac{{\mathrm{d}}a_{0}^{\mathrm{D}}}{{\mathrm{d}}u}$ are classical period and its dual, respectively.
For Chebyshev potentials, in the sector $u<1$, $\omega_{0}$ and $\omega^{\mathrm{D}}_{0}$ for the outermost well and the neighbouring barrier are Basar:2017hpr
$\omega_{0}(u)$ $=\frac{\sqrt{2}\pi}{m}\,\sin\left(\frac{\pi}{2m}\right)\,u\,{}_{2}F_{1}\left(\frac{1}{2}-\frac{1}{2m},\frac{1}{2}+\frac{1}{2m},1,u\right)\,,$ (94)
$\omega^{\mathrm{D}}_{0}(u)$ $=i\frac{\sqrt{2}\pi}{m}\,\sin\left(\frac{\pi}{m}\right)\,(1-u)\,{}_{2}F_{1}\left(\frac{1}{2}-\frac{1}{2m},\frac{1}{2}+\frac{1}{2m},2,1-u\right)\,.$ (95)
Then, the associated the classical action and their dual are written as
@@
$\omega^{\mathrm{D}}(u)=i\sqrt{\k{_}m}\,\omega(1-u)\,,\qquad a^{\mathrm{D}}(u)=-i\sqrt{\k{_}m}\,a(1-u)\,.$ (98)
A special case of the Chebyshev potentials is the periodic potential for which the classical action and its dual are written as
@@
[…]
$\omega^{\mathrm{D}}_{0}(u)=i\,\omega_{0}(1-u)\,,\qquad a^{\mathrm{D}}_{0}(u)=-i\,a_{0}(1-u)\,.$ (101)
@@
Let us start with the transition for the sector $u<1$ to the $u>1$ sector. Using the expressions in (94)-(97), it is straightforward to capture the evolution of (93) during the transition. Note that for $u>0$, $a_{0}^{\mathrm{D}}(u)(\omega^{\mathrm{D}}_{0}(u))$ is an analytic function while $a_{0}(u)(\omega_{0}(u))$ has a logarithmic branch-cut singularity at $u=1$ and to extend the expressions to $u>1$ sector, an analytic continuation is needed to be imposed.
To capture the effect of the branch cut of $a_{0}(u)$, we introduce $u\rightarrow u\pm i\varepsilon$ which is equivalent to the analytic continuation of the Stokes diagrams we discuss in Section 2.1. Then, using (NIST:DLMF, , Eqs. 15.2.3, 15.8.1): , we obtain the discontinuity on the branch cut as
@@
$\tau(u)=\frac{\omega_{0}^{\mathrm{D}}(u)}{\omega_{0}(u)},\qquad\tilde{\tau}(u_{\mathcal{D}})=\frac{\tilde{\omega}_{0}^{\mathrm{D}}(u_{\mathcal{D}})}{\tilde{\omega}_{0}(u_{\mathcal{D}})}\,,$ (107)
@@
$\omega_{0}^{\mathrm{D}}(u)=\sqrt{\k{_}m}\,\omega_{0}(u)\,,\qquad\tilde{\omega}_{0}^{\mathrm{D}}(u_{\mathcal{D}})=\frac{\tilde{\omega}_{0}^{\mathrm{D}}(u_{\mathcal{D}})}{\sqrt{\k{_}m}}\,.$ (108)
@@
$a_{0}^{\mathrm{D}}(u_{\mathcal{D}})=\sqrt{\k{_}m}\,a_{0}(u_{\mathcal{D}})\,,\qquad\tilde{a}_{0}^{\mathrm{D}}(u_{\mathcal{D}})=\frac{\tilde{a}_{0}^{\mathrm{D}}(u_{\mathcal{D}})}{\sqrt{\k{_}m}}\,.$ (109)
@@
$2i\,\tilde{a}^{\mathrm{D}}_{0}(0)\,\tilde{\omega}_{0}(0)=\tilde{a}_{0}(u_{\mathcal{D}})\,\tilde{\omega}^{\mathrm{D}}_{0}(u_{\mathcal{D}})-\tilde{\omega}_{0}(u_{\mathcal{D}})\,\tilde{a}_{0}^{\mathrm{D}}(u_{\mathcal{D}})\,,$ (110)
@@
maps the P-NP relation (92) for the potential $V$ to the one for the dual theory characterized by $V_{\mathrm{D}}=1-V$ and we conjecture
@@ 4 Periodic potential
In this section, we will apply the Airy and Weber-type EWKB methods to quantize one-dimensional periodic potential in all sectors and examine the connection between different sectors. Throughout this section, we consider the potential
@@
    $\Pi_{A}$ $=e^{-\frac{2\pi i}{\hbar}a(u,g)}=e^{-2\pi i\mathcal{F}\left(\frac{u}{g},g\right)}\,,$ (117)
    $\Pi_{B}$ $=e^{-\frac{2\pi i}{\hbar}a^{\mathrm{D}}(u,g)}=e^{-\mathcal{G}(\frac{u}{\hbar},g)}\frac{2\pi\left(\frac{g}{\mathcal{C}_{\mathrm{P}}}\right)^{-2\mathcal{F}\left(\frac{u}{\hbar},\hbar\right)}}{\left[\Gamma\left(\frac{1}{2}+\mathcal{F}\left(\frac{u}{g},g\right)\right)\right]^{2}}\,.$ (118)
@@
    $\Pi_{\tilde{A}}$ $=e^{-\frac{2\pi i}{\hbar}a(u,g)}=e^{-\mathcal{G}^{\mathrm{D}}(i\frac{1-u}{g},ig)}\frac{2\pi\left(\frac{-i\,g}{\mathcal{C}_{\mathrm{P}}}\right)^{-2\mathcal{F}^{\mathrm{D}}\left(i\frac{1-u}{g},ig\right)}}{\left[\Gamma\left(\frac{1}{2}+\mathcal{F}^{\mathrm{D}}\left(i\frac{1-u}{g},ig\right)\right)\right]^{2}}\,,$ (119)
    $\Pi_{B}$ $=e^{-\frac{2\pi i}{\hbar}a^{\mathrm{D}}(u,g)}=e^{-2\pi i\,\left(\mathcal{F}^{\mathrm{D}}\left(i\frac{1-u}{g},ig\right)\right)}\,.$ (120)
@@
where $\omega=1$ for $u<u_{\mathrm{T}}$ case and $\omega=i$ for $u>u_{\mathrm{T}}$ case. Note also that the integral of the constant $\mathcal{C}_{\mathrm{P}}$ around the barrier top is written in a slightly different form in view of (80). But it is straightforward to bring it to the form of (121). Finally, the functions $\mathcal{F}$, $\mathcal{G}$ and their dual $\mathcal{F}^{\mathrm{D}}$, $\mathcal{G}^{\mathrm{D}}$ in (117)-(120) can be expressed in series expansions of $u$ and $g$. We provide the first few orders at the end of this section, when we discuss the S-duality between $u\sim 0$ and $u\sim u_{\mathrm{T}}$ regions.
@@
The $u>u_{\mathrm{T}}$ sector, on the other hand, is less understood. Commonly, it was studied using ordinary WKB approximation Basar:2015xna . In Basar:2015xna , Başar and Dunne carried out a resurgence analysis considering all values of $u$. Of particular importance for our discussion, they analyzed the connection between the two sectors using the all order WKB method. They showed the smooth transition of the trans-series of $u$ between the two sectors despite the singularity of the WKB action $a(u)$ at $u=u_{\mathrm{T}}$ which should be taken care of via appropriate analytic continuations.
@@
  • We discuss the analytic continuation of the Stokes geometry between the two sectors based on our construction in Section 2.1. We show that the smooth transition of the spectrum shows itself in the median summation of the exact quantization conditions which stay intact upon the transition with appropriate redefinitions of the WKB cycles. In this way, we make the estimations in Basar:2015xna manifestly exact.

### v2

@@ 2.3 Dictionary
Let us finally discuss the connection between the Airy and Weber type EWKB in a more quantitative manner. In principal, the WKB cycles in the Weber-type approach are limits of the Airy-type ones at $u=u_{0}$, corresponding to minima or maxima of the potential. While the limit of cycles is not directly apparent in the EWKB framework, from a geometric point of view, it is a singular limit of the smooth (hyper)-elliptic curve corresponding to either $A$ or $B$ Airy-type cycle. Moreover, the singular curve is linked to the Euclidean instanton configurations, i.e bions or bounces, as we observed in the Weber-type EWKB. An explicit construction in these lines was made in Behtash:2017rqj , where the link between elliptic curves appearing in SUSY QM and complex-bion configuration was presented [^7] We thank Alireza Behtash for pointing out the geometric construction linking the singular limits of the WKB cycles and instanton configurations..
In the EWKB framework, the connection between Airy and Weber type cycles is provided by a dictionary. Note that this was originally discussed in Sueishi:2021xti ; DDP2 , but our construction covers more general cases. Moreover, stating the explicit dictionary serves as a guideline in later sections.
To make a direct connection to the literature which uses the all-order WKB expansion (e.g. see Basar:2017hpr ; Basar:2015xna ), let us recall the canonical definitions of the WKB action and its dual:
@@
Then, recalling the construction in Sueishi:2021xti and the relationship $u=g\tilde{u}$, we put forward a dictionary between Airy and Weber type cycles as follows:
@@
      $\Pi_{A}$ $=e^{-\frac{2\pi i}{g}a(u,g)}=e^{-\mathcal{G}^{\mathrm{D}}(i\frac{u}{g},ig)}\,\frac{2\pi\left(\frac{g}{\mathcal{C}_{\mathrm{i}}}\right)^{-2\mathcal{F}^{\mathrm{D}}(i\frac{u}{g},ig)}}{\left[\Gamma\left(\frac{1}{2}+\mathcal{F}^{\mathrm{D}}\left(i\frac{u}{g},ig\right)\right)\right]^{2}}\,,\quad\,(u_{0}=u_{\mathrm{T}})\,.$ (88)
@@
      $\Pi_{A}$ $=e^{-\frac{2\pi i}{g}a(u,g)}=e^{-\mathcal{G}^{\mathrm{D}}(i\frac{u}{g},ig)}\,\frac{\sqrt{2\pi}\left(\frac{g}{\mathcal{C}_{\mathrm{ii}}}\right)^{-\mathcal{F}^{\mathrm{D}}(i\frac{u}{g},ig)}}{\Gamma\left(\frac{1}{2}+\mathcal{F}^{\mathrm{D}}\left(i\frac{u}{g},ig\right)\right)}\,,\qquad(u_{0}=u_{\mathrm{T}})\,.$ (90)
@@
Note that we introduce $\mathcal{F}^{\mathrm{D}}$ and $\mathcal{G}^{\mathrm{D}}$ in the expressions (86), (88) and (90). They refer to the $S$-dual case characterized by $V_{\mathrm{D}}=u_{\mathrm{T}}-V$. The appearing imaginary factors are related to the inversion of a well or equivalently to $\omega_{0}=i$ case. This is the situation when we compute the functions $\mathcal{F}$ and $\mathcal{G}$ around a barrier top which are related to the dual functions as
@@ 3 Chebyshev potentials, transition and duality
    We elaborate on in their discussion with a focus on the transition from $u<1$ sector to $u>1$ sector, our findings in this section provide guidelines for the later sections where we analyze periodic and symmetric double-well potentials.
@@
    We also provide a conjecture on an $S$-duality transformation of the well-known P-NP relation Alvarez2 ; Alvarez3 ; Alvarez4 ; Dunne:2013ada ; Dunne:2014bca for genus-1 potentials in sector $u<1$.
@@
A resurgence relation between perturbative and non-perturbative physics can be expressed by the WKB actions $a(u,g)$ and $a^{\mathrm{D}}(u,g)$ or equivalently $\mathcal{F}(\tilde{u},g)$ and $\mathcal{G}(\tilde{u},g)$. In the EWKB framework, it is more natural to obtain a relation between $\mathcal{F}$ and $\mathcal{G}$: Around the bottom of wells and for genus-1 potentials [^9] More specifically, (92) is for the cases when all the actions of the wells are proportional to each others which indicates there are only two independent WKB actions. Recently, a generalization was presented in Cavusoglu:2023bai ; Cavusoglu:2024usn . The authors showed a paramteric form of P-NP holds for the deformed quartic potential which reduces to cubic and quartic cases in extreme limits and P-NP relation takes the form of (92)., this relationship is well-known in various forms Alvarez2 ; Alvarez3 ; Alvarez4 ; Dunne:2013ada ; Dunne:2014bca and in general, it can be expressed as
@@
$\mathsf{w}_{0}^{\mathrm{D}}(u)\,a_{0}(u)-a_{0}^{\mathrm{D}}(u)\,\mathsf{w}_{0}(u)=2ia_{0}^{\mathrm{D}}(0)\,\mathsf{w}_{0}(0)\,,$ (93)
where $\mathsf{w}_{0}=\frac{{\mathrm{d}}a_{0}}{{\mathrm{d}}u}$ and $\mathsf{w}_{0}^{\mathrm{D}}=\frac{{\mathrm{d}}a_{0}^{\mathrm{D}}}{{\mathrm{d}}u}$ are classical period and its dual, respectively.
For Chebyshev potentials, in the sector $u<1$, $\mathsf{w}_{0}$ and $\mathsf{w}^{\mathrm{D}}_{0}$ for the outermost well and the neighbouring barrier are Basar:2017hpr
$\mathsf{w}_{0}(u)$ $=\frac{\sqrt{2}\pi}{m}\,\sin\left(\frac{\pi}{2m}\right)\,u\,{}_{2}F_{1}\left(\frac{1}{2}-\frac{1}{2m},\frac{1}{2}+\frac{1}{2m},1,u\right)\,,$ (94)
$\mathsf{w}^{\mathrm{D}}_{0}(u)$ $=i\frac{\sqrt{2}\pi}{m}\,\sin\left(\frac{\pi}{m}\right)\,(1-u)\,{}_{2}F_{1}\left(\frac{1}{2}-\frac{1}{2m},\frac{1}{2}+\frac{1}{2m},2,1-u\right)\,.$ (95)
Then, the associated classical action and their dual are written as
@@
$\mathsf{w}^{\mathrm{D}}(u)=i\sqrt{\k{_}m}\,\mathsf{w}(1-u)\,,\qquad a^{\mathrm{D}}(u)=-i\sqrt{\k{_}m}\,a(1-u)\,.$ (98)
Finally, note that periodic potential is a special case of the Chebyshev potentials. In this case, the classical action and its dual are written as
@@
and the period(action) and its dual are related to each other by
$\mathsf{w}^{\mathrm{D}}_{0}(u)=i\,\mathsf{w}_{0}(1-u)\,,\qquad a^{\mathrm{D}}_{0}(u)=-i\,a_{0}(1-u)\,.$ (101)
which corresponds to the case of $\k{_}\mathrm{P}=1_$ in comparison with (98).
@@
Let us start with the transition from the sector $u<1$ to the $u>1$ sector. Using the expressions in (94)-(97), it is straightforward to capture the evolution of (93) during the transition. Note that for $u>0$, $a_{0}^{\mathrm{D}}(u)(\mathsf{w}^{\mathrm{D}}_{0}(u))$ is an analytic function while $a_{0}(u)(\mathsf{w}_{0}(u))$ has a branch-cut along the real axis for $u\geq 1$. Then, in the $u>1$ sector, an analytic continuation is needed to be imposed for $a_{0}(u)(\mathsf{w}_{0}(u))$. For this reason, we introduce $u\rightarrow u\pm i\varepsilon$. Note that this is equivalent to the analytic continuation of the Stokes diagrams for phases $\theta_{u}=\pm\pi^{\mp}$ we discuss in Section 2.1.
As $a_{0}$ and $a_{0}^{\mathrm{D}}$ are given by the hypergeometric function, we use its analytic properties (NIST:DLMF, , Eqs. 15.2.3, 15.8.1) to obtain the discontinuity on the branch cut as
@@
$\tau(u)=\frac{\mathsf{w}_{0}^{\mathrm{D}}(u)}{\mathsf{w}_{0}(u)},\qquad\tilde{\tau}(u_{\mathcal{D}})=\frac{\tilde{\mathsf{w}}_{0}^{\mathrm{D}}(u_{\mathcal{D}})}{\tilde{\mathsf{w}}_{0}(u_{\mathcal{D}})}\,,$ (107)
@@
$\tilde{\mathsf{w}}_{0}(u_{\mathcal{D}})=\sqrt{\k{_}m}\,\mathsf{w}_{0}(u_{\mathcal{D}})\,,\qquad\tilde{\mathsf{w}}_{0}^{\mathrm{D}}(u_{\mathcal{D}})=\frac{\mathsf{w}_{0}^{\mathrm{D}}(u_{\mathcal{D}})}{\sqrt{\k{_}m}}\,.$ (108)
@@
$\tilde{a}_{0}(u_{\mathcal{D}})=\sqrt{\k{_}m}\,a_{0}(u_{\mathcal{D}})\,,\qquad\tilde{a}_{0}^{\mathrm{D}}(u_{\mathcal{D}})=\frac{a_{0}^{\mathrm{D}}(u_{\mathcal{D}})}{\sqrt{\k{_}m}}\,.$ (109)
@@
$2i\,\tilde{a}^{\mathrm{D}}_{0}(0)\,\tilde{\mathsf{w}}_{0}(0)=\tilde{a}_{0}(u_{\mathcal{D}})\,\tilde{\mathsf{w}}^{\mathrm{D}}_{0}(u_{\mathcal{D}})-\tilde{\mathsf{w}}_{0}(u_{\mathcal{D}})\,\tilde{a}_{0}^{\mathrm{D}}(u_{\mathcal{D}})\,,$ (110)
@@
maps the P-NP relation (92) for the potential $V$ to the one for the dual theory characterized by $V_{\mathrm{D}}=1-V$ and conjecture
@@ 4 Periodic potential
In this section, we apply the Airy and Weber-type EWKB methods to quantize one-dimensional periodic potential in all sectors and examine the connection between different sectors. Throughout this section, we consider the potential
@@
    $\Pi_{B}$ $=e^{-\frac{2\pi i}{g}a^{\mathrm{D}}(u,g)}=e^{-\mathcal{G}(\frac{u}{g},g)}\frac{2\pi\left(\frac{g}{\mathcal{C}_{\mathrm{P}}}\right)^{-2\mathcal{F}\left(\frac{u}{g},g\right)}}{\left[\Gamma\left(\frac{1}{2}+\mathcal{F}\left(\frac{u}{g},g\right)\right)\right]^{2}}\,.$ (118)
@@
    $\Pi_{\tilde{A}}$ $=e^{-\frac{2\pi i}{g}a(u,g)}=e^{-\mathcal{G}^{\mathrm{D}}(i\frac{1-u}{g},ig)}\frac{2\pi\left(\frac{g}{\mathcal{C}_{\mathrm{P}}}\right)^{-2\mathcal{F}^{\mathrm{D}}\left(i\frac{1-u}{g},ig\right)}}{\left[\Gamma\left(\frac{1}{2}+\mathcal{F}^{\mathrm{D}}\left(i\frac{1-u}{g},ig\right)\right)\right]^{2}}\,,$ (119)
    $\Pi_{B}$ $=e^{-\frac{2\pi i}{g}a^{\mathrm{D}}(u,g)}=e^{-2\pi i\,\left(\mathcal{F}^{\mathrm{D}}\left(i\frac{1-u}{g},ig\right)\right)}\,.$ (120)
@@
where $\omega=1$ for $u<u_{\mathrm{T}}$ case and $\omega=i$ for $u>u_{\mathrm{T}}$ case. Note also that initially the integral of the constant $\mathcal{C}_{\mathrm{P}}$ around the barrier top is written in a slightly different form in view of (80). But it is straightforward to bring it to the form of (121). Finally, the functions $\mathcal{F}$, $\mathcal{G}$ and their dual $\mathcal{F}^{\mathrm{D}}$, $\mathcal{G}^{\mathrm{D}}$ in (117)-(120) can be expressed in series expansions of $u$ and $g$. We provide the first few orders at the end of this section, when we discuss the $S$-duality between $u\sim 0$ and $u\sim u_{\mathrm{T}}$ regions.
@@
The $u>u_{\mathrm{T}}$ sector, on the other hand, is less studied Connor84 ; weinstein1985hill ; weinstein1987asymptotic ; Basar:2015xna . Of particular importance for our discussion, in Basar:2015xna , Başar and Dunne carried out a resurgence analysis using all order WKB expansions for all values of $u$. They analyzed the connection between the two sectors using the all order WKB method. They showed the smooth transition of the trans-series of $u$ between the two sectors despite the singularity of the WKB action $a(u)$ at $u=u_{\mathrm{T}}$ which should be taken care of via appropriate analytic continuations.
@@
  • We discuss the analytic continuation of the Stokes geometry between the two sectors based on our construction in Section 2.1. We show that the smooth transition of the spectrum shows itself in the median summations of the exact quantization conditions which stay intact upon the transition with appropriate redefinitions of the WKB cycles. In this way, we make the estimations in Basar:2015xna manifestly exact.